OpenStudy (anonymous):

how do you solve this problem? :o 2+[3(-4)-4]=2(-4-3)

OpenStudy (beefarm):

The most important thing here is to understand the order of operations, which go like this: 1. Parentheses 2. Exponents 3. Multiplication and Division 4. Addition and Subtraction The steps for solving this particular problem are as follows: 2+[3*(-4)-4]=2*(-4-3) //I have just added * where multiplication occurs, just for clarity 2+[-12-4]=2*(-4-3) //The first part to evaluate is the multiplication that occurs inside a set of parentheses 2+(-16)=2*(-7) //Next, finish evaluating the expressions within parentheses 2+(-16)=-14 //Next look for exponents, there are none here, look for multiplication or division, there is a multiplication so we evaluate it next -14=-14 //Next do any remaining addition and subtraction Hope that helped, good luck! Let me know if any of that doesn't make sense, haha

OpenStudy (anonymous):

Thank You so much!!! :D